OpenStudy (anonymous):

solve by completing the square. x squared -2x=3. show work please.

OpenStudy (lgbasallote):

x^2 -2x = 3 x^2 -2x + (2/2)^2 = 3 + (2/2)^2 x^2 - 2x + 1 = 3 + 1 (x-2)^2 = 4 x-2 = sqrt (4) x-2 = +/- 2 x = 2 +/- 2 x = 4 x=0

OpenStudy (ash2326):

@ineedhelppleaseee did you understand?

OpenStudy (anonymous):

completely :) for this promblem